OpenStudy (anonymous):

Can someone answer this for me please? Which three ordered pairs complete the table to give solutions of the equation y = x – 4? A. (–3, –7), (0, –4), and (3, –1) B. (–3, –7), (0, –4), and (3, 1) C. (–3, –1), (0, –4), and (3, 1) D. (–3, –1), (0, –4), and (3, –1)

OpenStudy (anonymous):

Hi Welcome to Openstudy! I would suggest that you should try plugging in the x-coordinates of the points into the given equation and the set of points that will satisfy the equation will be the answer. For example, let's try choice A. It has a point (-3,-7). x=-3 plug it in: y = x – 4 y=-3-4 y=-7 so this point satisfy the equation. Do this on the next point and the rest of the choices.

OpenStudy (anonymous):

k thanks